OpenStudy (anonymous):

solve cos2x(2cosx+1)=0

OpenStudy (anonymous):

2cos(x)cos(2x)+cos(2x)=0

OpenStudy (anonymous):

2cos(a)cos(b)=cos(a+b)+cos(a-b)

OpenStudy (anonymous):

Just keep the equation as it is. \[\cos(2x)(2\cos(x)+1)=0\]This means that either \[\cos(2x)=0\] or \[\ 2cos(x)+1=0\] Ought to be fairly simple onwards.

OpenStudy (lgbasallote):

\[\cos^2 x(2\cos x + 1) = 0?\]

OpenStudy (anonymous):

nice work....@Uniquebum

OpenStudy (anonymous):

cos(2x)=0 \[2x=n*\pi/2 \implies x=n*\pi/4\]

OpenStudy (anonymous):

n=1,2,3..

OpenStudy (anonymous):

2cosx+1=0 2cosx=-1 cosx=-1/2